OpenStudy (anonymous):

the limit of f(x) exists at x = a, which statement is true? limx→a−f(x)=limx→a+f(x)=L limx→af(x)=limx→a+f(x)=L limx→a−f(x)=limx→af(x)=L limx→a−f(x)=L

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

@tkhunny

OpenStudy (michele_laino):

please can you rewrite your exercise, using the editor of formulas?

OpenStudy (anonymous):

There is also a x with a arrow and a A^-

OpenStudy (anonymous):

I will try my best to draw

OpenStudy (perl):

I am having trouble reading that too .

OpenStudy (michele_laino):

I can say this: if: \[\mathop {\lim }\limits_{x \to a} f\left( x \right) = L\] then: we can show, that \[\mathop {\lim }\limits_{x \to a} \left( { - f\left( x \right)} \right) = - L\] that is a theorem of the theory

OpenStudy (anonymous):



OpenStudy (anonymous):

triedd my best

OpenStudy (michele_laino):

ok! I understand now

OpenStudy (anonymous):

That was A

OpenStudy (michele_laino):

By definition of the limit of a function, if exist the subsequent limit: \[\mathop {\lim }\limits_{x \to a} f\left( x \right) = L\] then also the right limit and the left limit exist That is a theorem of the Mathematical Analysis

OpenStudy (michele_laino):

So, what is the right option?

OpenStudy (anonymous):

hold up Im lagging

OpenStudy (perl):

a) looks right, what about the other options

OpenStudy (anonymous):

yea A

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thx yall

OpenStudy (michele_laino):

thank you! @DaWizjr @perl

OpenStudy (perl):

all the choices look correct